(((Sove 3x^2 -2x +4 = 0 by the quadratic Formula)))
Answer by ikdeep(226)

3x^2 -2x +4 = 0
since here we can't split the middle term so we have to find DISCRIMINANT
d OR DISCRIMINANT = b^2 - 4ac
in the above equation a = 3 , b = -2 and c = 4
put the values of a , b nad c in the above formula and we get
d = 4 - 48
d = -44
since d < 0 then the above equation have imaginary roots or we can say it dnt have roots
Hope this will help you
Please feel free to revert back for any further queries.